OpenStudy (desmarie):

dilute 10.0 mL of the 0.50 M NaCl by adding distilled water until 100 mL of solution are produced. Does the 100 mL of diluted NaCl have more, less, or the same quantity of NaCl as the original 10.0 mL of 0.50 M NaCl?

OpenStudy (desmarie):

@SLP-HUERTA @Photon336 @Cuanchi

OpenStudy (desmarie):

would the answer be that they have the same and if yes why?

OpenStudy (desmarie):

@Tennis5518

OpenStudy (desmarie):

haha

OpenStudy (tennis5518):

Any clues?

OpenStudy (tennis5518):

C1 = original concentration V1 = original volume C2 = final concentration V2 = final volume C1 x V1 = C2 x V2 C2 = (C1 x V1)/C2 C2 = (0.5 x 10)/100 C2 = 0.05

OpenStudy (desmarie):

@Photon336

OpenStudy (photon336):

So we're diluting 10.0 mL of the 0.50 M NaCl by adding distilled water until 100 mL of the solution is produced. well, I guess we can figure out the new concentration. \[N_{a}V_{a} = N_{b}V_{B}\]

OpenStudy (photon336):

Na = 0.50M Va = 10mL Nb = ? Vb = 100mL

OpenStudy (tennis5518):

It has the same quantity. You are not adding NaCl, and it doesn't just vanish. All that changes is the concentration of NaCl (the amount per unit volume), but the amount is constant.

OpenStudy (tennis5518):

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Tennis5518 C1 = original concentration V1 = original volume C2 = final concentration V2 = final volume C1 x V1 = C2 x V2 C2 = (C1 x V1)/C2 C2 = (0.5 x 10)/100 C2 = 0.05 \(\color{#0cbb34}{\text{End of Quote}}\) It true

OpenStudy (photon336):

\[\frac{ NaVa }{ Vb } = Nb \] \[\frac{ 0.50M*0.01\cancel\L }{ 0.1\cancel\L } = 0.05M\]

OpenStudy (photon336):

yeah I agree with @Tennis5518 like you have the same amount of NaCl it's just that it's concentration is not as much.

OpenStudy (desmarie):

ok thanks so much for your help :)

OpenStudy (tennis5518):

Np :)